OpenStudy (madeleine2):

Write the standard form of the line that passes through the point (-2, 4) and is parallel to x - 2y = 6

563blackghost (563blackghost):

We need to change the standard form equation into slope intercept form. \(\huge\bf{x−2y=6}\) Subtract x from both sides. \(\huge\bf{−2y=-x + 6}\) Next divide by -2 on each side. \(\huge\bf{y=\frac{1}{2}x−3}\) ~~~~~~~~~~~~ Now parallel lines contain the same slope but different y-intercepts so we keep the slope. \(\huge\bf{y=\frac{1}{2}x+b}\) To find b we would input the point we want it to go through. \(\huge\bf{4=\frac{1}{2}(−2)+b}\) Simplify. What does b equal?

OpenStudy (madeleine2):

-3? @563blackghost

OpenStudy (steve816):

Close! Can you try again? \[\huge 4=-1+b\]

OpenStudy (madeleine2):

5? @steve816

OpenStudy (steve816):

Coorect! NIce job :)